find an equation in slope intercept form for the line.
through (-2,3) perpendicular to y=2x+5

Answer by TutorDelphia(193)

Perpendicular lines have slopes that are the opposite reciprocal (flip it over and make it negative).
Is our slope is 2/1 so the new slope will be -1/2

So far we know
y=(-1/2)x+b
We can sub in -2 and 3 for x and y and solve for b
3=(-1/2)*(-2)+b
3=1+b
b=2
now going back to y=mx+b
we know that m=(-1/2) and b=2
y=(-1/2)x+2 is our perpendicular line